Describe and correct the error in finding the probability of randomly drawing the given card from a standard deck of 52 cards.
p(heart or face card)=p(heart)+p(face card) = 13/52+12/52=25/52.
So it's asking me to find what's wrong with the way they solved this but I don't see any mistake.

Answer by richard1234(7193)

With the way you solved it, you overcounted. In other words, you counted the face card hearts twice (jack, queen, king of hearts). We subtract the probability of obtaining a face card of the heart suit, which is 3/52, to obtain 25/52 - 3/52 = 22/52 = 11/26.